OpenStudy (anonymous):

F : x > 3 - x what is the value of F(7)

OpenStudy (anonymous):

f(7)=3-7=-4 replace x with 7

OpenStudy (anonymous):

7>-4 and thus 11>0

OpenStudy (anonymous):

lost me on the last one

OpenStudy (anonymous):

i am guessing that the the function notation is \[F: x\mapsto 3-x\] and therefore as you said \[x\mapsto 3-7=-4\] but then \[11\mapsto 3-11=-7\] right?